OpenStudy (anonymous):

A family has 8 girls and 4 boys. A total of 2 children must be chosen to speak on the behalf of the family at a local benefit. What is the probability that 1 girl and 1 boy will be chosen?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

8 ways to choose a girl 4 ways to choose a boy and \(\binom{12}{2}=\frac{12\times 11}{2}=6\times 11=66\) ways to choose two out of 12

OpenStudy (misty1212):

so \[\frac{8\times 4}{66}\] is your probability

OpenStudy (anonymous):

Thank you so much, I really appreciate your help

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]